OpenStudy (anonymous):

What is the value of the expression 2^2 + 6^2 ÷ 2^2? 10 13 16 36

OpenStudy (anonymous):

@peachpi @taylorz115820

OpenStudy (anonymous):

its 13 right?

jagr2713 (jagr2713):

Correct but do you still need help on it?

OpenStudy (anonymous):

I dont think so...

OpenStudy (anonymous):

Jar how did you get 13?

jagr2713 (jagr2713):

Oh ok :D @taylorz115820 what would we do first if you were answering this?

Mehek (mehek14):

\(36/4 = 9 \\ 9+4 = 13 \) @taylorz115820

OpenStudy (anonymous):

OMG I forgot about order of operations my bad

jagr2713 (jagr2713):

Yea :D PEMDAS

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

if you didn't do that you would have gotten 10 glad someone else answered and not me that would have been bad

jagr2713 (jagr2713):

Its ok :D Now you know :D

OpenStudy (mathmate):

Is it A. \(2^2+3^2\div 2^2\), or B. \(\Large \frac{2^2+3^2}{2^2}\), or